OpenStudy (peterg2001):

help please file:///home/chronos/u-f52605ba6a4fa8f1c54eef593e9591f629bc4268/Downloads/Pre-Algebra%202%20Unit%202%20Summary%20Assignment%20(1).pdf

OpenStudy (alexandervonhumboldt2):

it doesnt open

OpenStudy (anonymous):

hes right

OpenStudy (peterg2001):

\[Question 1: Solve- 3y+6\le -15 and graph the solution on a number line.\]

OpenStudy (peterg2001):



OpenStudy (alexandervonhumboldt2):

-3y<=-21 y<=7

OpenStudy (peterg2001):

Question 2: Evaluate 3/21divide 6/7 and write in lowest terms.

OpenStudy (peterg2001):

Question 3: Evaluate -3 5/8+7 3/4 and write as a mixed number.

OpenStudy (peterg2001):

done

OpenStudy (peterg2001):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

2. 3/21*7/6=21/126=1/6

OpenStudy (alexandervonhumboldt2):

3. -3 5/8+7 3/4=4 11/8=5 3/8

OpenStudy (peterg2001):

In order to receive credit, you must SHOW ALL YOUR WORK that is necessary to answer the question

OpenStudy (peterg2001):

@Directrix

OpenStudy (peterg2001):

@hhelpplzzzz

OpenStudy (peterg2001):

n order to receive credit, you must SHOW ALL YOUR WORK that is necessary to answer the question @AlexandervonHumboldt2

OpenStudy (peterg2001):

@geneticrockhopper247